# XiHe: A Data-Driven Model for Global Ocean Eddy-Resolving Forecasting

### v2

@@ Abstract
Global ocean forecasting is fundamentally important to support marine activities. The leading operational Global Ocean Forecasting Systems (GOFSs) use physics-driven numerical forecasting models that solve the partial differential equations with expensive computation. Recently, specifically in atmosphere weather forecasting, data-driven models have demonstrated significant potential for speeding up environmental forecasting by orders of magnitude, but there is still no data-driven GOFS that matches the forecasting accuracy of the numerical GOFSs. In this paper, we propose the first data-driven $1/12^{\circ}$ resolution global ocean eddy-resolving forecasting model named XiHe, which is established from the 25-year France Mercator Ocean International’s daily GLORYS12 reanalysis data. XiHe is a hierarchical transformer-based framework coupled with two special designs. One is the land-ocean mask mechanism for focusing exclusively on the global ocean circulation. The other is the ocean-specific block for effectively capturing both local ocean information and global teleconnection. Extensive experiments are conducted under satellite observations, in situ observations, and the IV-TT Class 4 evaluation framework of the world’s leading operational GOFSs from January 2019 to December 2020. The results demonstrate that XiHe achieves stronger forecast performance in all testing variables than existing leading operational numerical GOFSs including Mercator Ocean Physical SYstem (PSY4), Global Ice Ocean Prediction System (GIOPS), BLUElinK OceanMAPS (BLK), and Forecast Ocean Assimilation Model (FOAM). Particularly, the accuracy of ocean current forecasting of XiHe out to 60 days is even better than that of PSY4 in just 10 days. Additionally, XiHe is able to forecast the large-scale circulation and the mesoscale eddies. Furthermore, it can make a 10-day forecast in only 0.36 seconds, which accelerates the forecast speed by thousands of times compared to the traditional numerical GOFSs.
@@ 1 Introduction
Ocean forecasting is critically important for many marine activities. At present, the leading GOFSs (e.g. Mercator Ocean Physical SYstem (PSY4) and Real-Time Ocean Forecast System (RTOFS)) use physics-driven models in fluid mechanics and thermodynamics to predict future ocean motion states and phenomena based on current ocean conditions [1]. The GOFSs adopt numerical methods that rely on supercomputers to solve the partial differential equations of the physical models. Due to their desirable performance, they are operationally run in different countries worldwide. However, numerical forecasting methods are usually computationally expensive and slow. For example, a single forecasting simulation in the numerical GOFSs may take hours on a supercomputer with hundreds of computational nodes [2]. Besides, improving the forecasting accuracy of these methods is exceedingly challenging because they heavily rely on the human cognitive abilities in understanding the physical laws of the ocean environment [3].
@@
    We propose the first data-driven $1/12^{\circ}$ resolution global ocean eddy-resolving forecasting model XiHe, which achieves stronger forecast performance in all testing variables than the compared leading operational numerical GOFSs. Moreover, XiHe can make a forecast in only 0.36 seconds (1000x faster than the numerical GOFSs) averagely with one GPU.
@@ 2.1 Physics-Driven Methods
Until now, ocean numerical forecasting methods have contributed overall the highest prediction accuracy, but they still face two major limitations. Firstly, they are usually computationally expensive and slow due to the high complexity of solving the physical partial differential equations [4]. Secondly, the improvement of the forecasting accuracy is exceedingly challenging due to their limitation of human cognitive abilities in understanding the physical laws of the ocean environment [3].
@@ 3.1 Training Dataset
Considering the impact of the sea surface wind on ocean dynamics, we incorporate the 10m u-component of wind (U10) and 10m v-component of wind (V10) into the training data of XiHe. The wind field data for training is from the Fifth Generation Global Atmospheric Reanalysis data (ERA5) produced by the European Centre for Medium-Range Weather Forecasts (ECMWF) [30]. Their spatial resolution is $0.25^{\circ}$ with 137 model levels. In order to standardize the data dimensions, we interpolated the wind data to the same spatial and temporal resolution as the GLORYS12 reanalysis data.
@@ 4.4.2 Global SIE module
${\mathbf{G}^{\prime}}^{l}=\operatorname{Concat}_{\{h\}}\left(\operatorname{Attention}\left(\mathbf{W}_{h}^{Q}\mathbf{G}_{h}^{l},\mathbf{G}_{h}^{K}\mathbf{\tilde{Z}}^{l}_{h},\mathbf{W}_{h}^{V}\mathbf{\tilde{Z}}^{l}_{h}\right)\right),$ (5)
@@ 5 Experimental Evaluation
The forecasting capacity of the XiHe is evaluated based on satellite observations, in situ observations, and the IV-TT Class 4 evaluation framework of the world’s leading operational GOFSs. Note that XiHe is trained on 25 years of data (from 1993 to 2017), validated on data of 2018, and tested on data of 2019 and 2020. It forecasts the oceanic variables in a $1/12^{\circ}$ spatial resolution (grid format of 2041$\times$4320), which is the same as the GLORYS12 reanalysis data.
@@ 5.1.1 The IV-TT Class 4 Framework
We evaluated the performances of XiHe under the GODAE OceanView IV-TT Class 4 framework, which is designed to evaluate the world’s leading operational ocean forecasting systems like FOAM, PSY4, GIOPS, BLK, and RTOFS [32]. The IV-TT Class 4 framework focuses on the assessment of the forecast performance and skill, using available ocean observations including in situ drifting buoys and satellite data in near real-time. Within the IV-TT Class 4 framework, each forecasting system interpolates its forecast fields, including ocean temperature profile, ocean salinity profile, SLA, etc., to the observation locations. Then different forecasting systems can be evaluated using the observations and compared with each other. The framework utilizes ocean Temperature and Salinity observations that are sourced from Argo profiles. Additionally, SLA was measured by satellite altimeters [32].
There are five statistical metrics in the IV-TT Class 4 framework, including Bias, root mean square error (RMSE), ACOR, and two skill scores constructed using persistence and climatology fields [32]. The RMSE, ACOR, and Bias are used in this paper. According to the evaluation of the IV-TT framework, we first interpolate the grid forecast results into the IV-TT observations and then calculate the evaluation metrics [32]. For example, the RMSE and ACOR are defined in Eq. (10) and Eq. (11).
@@
${\rm ACOR}=\frac{\sum_{i=1}^{N}(\mathbf{F_{i}}-\mathbf{C_{i}})(\mathbf{O_{i}}-\mathbf{C_{i}})}{\sqrt{\sum_{i=1}^{N}(\mathbf{F_{i}}-\mathbf{C_{i}})^{2}}\sqrt{\sum_{i=1}^{N}(\mathbf{O_{i}}-\mathbf{C_{i}})^{2}}},$ (11)
@@
We compare XiHe with the four state-of-the-art global ocean forecasting systems, including PSY4 [17], GIOPS [15], FOAM [16], and BLK [14]. Details of these forecasting systems are listed in Table I. For RTOFS, considering that there are only four months of forecasting data from Jan. 2019 to Dec. 2020, thus we do not show the experimental results of RTOFS in this paper.
@@ 5.1.2 Overall Evaluation
For all the evaluated variables including ocean temperature, salinity, 15m-depth current, and SLA, XiHe consistently outperforms all the numerical GOFSs. The advantage persists across lead times from 1 day to 10 days. For ocean temperature, compared to FOAM, PSY4, GIOPS, and BLK in the 144-hour (6-day) lead time forecasting, XiHe’s vertical averaged temperature provides an average improvement of 2.51%, 3.91%, 3.78%, and 19.05%, respectively. Meanwhile, XiHe’s vertical averaged salinity provides an average improvement of 2.20%, 6.33%, 9.21%, and 25.21%, respectively.
@@
Compared with the PSY4 system on the 15m-depth zonal and meridional component of ocean currents, XiHe shows a much smaller growing tendency in RMSE. The slopes of the error curves can reflect the growth of forecasting error as the lead time increases. We calculated the approximate slopes of the error curves for PSY4 and XiHe with the lead times from 1 day to 10 days, respectively. For the 15m-depth zonal (meridional) component of ocean currents, the slope of PSY4 is 0.00189 (0.00180) $m/(s\cdot day)$, which is about 3.0 times (4.5 times) as large as 0.00064 (0.00040) $m/(s\cdot day)$ of XiHe. This demonstrates that the forecasting error of XiHe grows much slower than PSY4, and XiHe has the potential for a longer forecast lead time. It is noteworthy that the 10-day RMSE of XiHe is remarkably smaller than the 1-day RMSE of PSY4. Furthermore, to assess the forecasting performance of XiHe for a longer lead time, we train the models for the lead time of 20 days, 30 days, and 60 days. As shown in Fig 3, the RMSE values of the 15m-depth zonal component of ocean currents (uo) for the three lead times are 0.2041, 0.2093, and 0.2165 respectively, which is even smaller than the PSY4’s 10-day RMSE of 0.2214. The RMSE values of the 15m-depth meridional component of ocean currents (vo) for the three lead times are 0.1955, 0.2008, and 0.2062 respectively, which is smaller than the 10-day RMSE of 0.2149 for PSY4. This indicates that the accuracy of ocean current forecasting of XiHe out to 60 days is even better than that of PSY4 in 10 days.
@@
From the perspective of both the oceanic variables and the lead times, the overall evaluation results firmly demonstrate that the data-driven XiHe model has achieved stronger forecast performance than the leading operational numerical GOFSs including PSY4, GIOPS, BLK, and FOAM.
@@
Figure 4: The performance comparison of our data-driven model XiHe with the four leading operational numerical GOFSs based on the IV-TT framework. The x-axis represents the forecast time. The y-axis represents the forecasting ACOR (higher is better) of the global ocean forecasting systems. In addition, for each lead time, the ACOR value is averaged from Jan. 1st, 2019 to Dec. 31st, 2020.
@@
As shown in Fig. 4 of the ACOR analysis, XiHe outperforms the four leading operational numerical GOFSs in terms of temperature and salinity. The ACOR values of temperature and salinity profiles of XiHe’s 20-day forecasting are 0.6177 and 0.5260, which are even better than the 10-day forecasts of other GOFSs. This highlights XiHe’s robust forecasting capabilities.
@@
To be more detailed, we present the daily RMSE of temperature, salinity, uo, vo, and SLA from Jan. 1st, 2019 to Dec. 31st, 2020 in Fig. 6 with the lead times of 1-day, 5-day, and 10-day. For the five variables, XiHe exceeds the numerical forecasting systems by an immense margin during the evaluated period. Especially for the uo and vo variables, XiHe achieves an average improvement of 9.95% and 11.16% over PSY4 in the 10-day lead time forecasting, respectively. It can be observed that the forecast errors for different variables exhibit certain seasonal variation characteristics. For example, with regard to the XiHe and PSY4 models, the forecast errors for ocean currents (uo, vo) are smaller in the boreal summer and larger in the boreal winter. As the forecast lead time increases, the seasonal variation in forecast errors becomes more evident. In addition, it is worth mentioning that the forecast results provided by XiHe are more stable than PSY4, exhibiting fewer excessive peak values that would indicate dramatic errors.
@@ 5.1.3 Temperature/Salinity Profiles Evaluation
In order to gain more insights into the forecast errors at different depths, Fig. 7 and Fig. 8 show the profiles of the temperature and the salinity RMSE for different GOFSs in depth above 644.0m.
As to the depth-wise evaluation results towards the salinity variable in Fig. 7, XiHe significantly outperforms all the compared numerical operational GOFSs in the whole depths. This denotes that XiHe performs notably better than the leading GOFSs in accurately forecasting the salinity variable throughout the profile.
As shown in Fig. 3, XiHe gets the best performance in overall temperature evaluation. The experimental results in Fig. 8 show the depth-wise evaluation of temperature. XiHe consistently outperforms the four operational numerical GOFSs in terms of temperature variables almost in the whole depths, which is particularly excellent below the depth of 100m. For the temperature profiles above 100m, XiHe’s results don’t achieve significant improvements over the numerical GOFSs. The forecast error of XiHe is slightly larger than that of FOAM and GIOPS, comparable to BLK, and better than PSY4. The reason may be that the forecasting capacity of XiHe for shallow temperature is limited by the GLORYS12 reanalysis data. This is supported by the fact that PSY4, which also comes from the same forecasting system as the GLORYS12 reanalysis data, performs poorly in shallow temperature forecasting as well. However, despite this limitation, XiHe still significantly outperforms PSY4 and shows potential for stronger forecasting capacity with quality improvements of the reanalysis data for training in the future.
Figure 7: The profiles of salinity RMSE above 644m for different GOFSs. The lead times range from 1 day to 10 days. For each lead time, the reported RMSE is averaged from Jan. 1st, 2019 to Dec. 31st, 2020.
Figure 8: The profiles of temperature RMSE above 644m for different GOFSs. The lead times range from 1 day to 10 days. For each lead time, the reported RMSE is averaged from Jan. 1st, 2019 to Dec. 31st, 2020.

### v4

@@ Abstract
Global ocean forecasting is fundamentally important to support marine activities. The leading operational Global Ocean Forecasting Systems (GOFSs) use physics-driven numerical forecasting models that solve the partial differential equations with expensive computation. Recently, specifically in atmosphere weather forecasting, data-driven models have demonstrated significant potential for speeding up environmental forecasting by orders of magnitude, but there is still no data-driven GOFS that matches the forecasting accuracy of the numerical GOFSs. In this paper, we propose the first data-driven $1/12^{\circ}$ resolution global ocean eddy-resolving forecasting model named XiHe, which is established from the 25-year France Mercator Ocean International’s daily GLORYS12 reanalysis data. XiHe is a hierarchical transformer-based framework coupled with two special setups. One is the land-ocean mask mechanism for focusing exclusively on the global ocean circulation. The other is the ocean-specific block for effectively capturing both local ocean information and global teleconnection. Extensive experiments are conducted under satellite observations, in situ observations, and the IV-TT Class 4 evaluation framework of the world’s leading operational GOFSs from January 2019 to December 2020. The results demonstrate that XiHe achieves stronger forecast performance in all testing variables than existing leading operational numerical GOFSs including Mercator Ocean Physical SYstem (PSY4), Global Ice Ocean Prediction System (GIOPS), BLUElink OceanMAPS, and Forecast Ocean Assimilation Model (FOAM). Particularly, the accuracy of ocean current forecasting of XiHe out to 60 days is even better than that of PSY4 in just 10 days. Additionally, XiHe is able to forecast the large-scale circulation and the mesoscale eddies. Furthermore, it can make a 10-day forecast in only 0.35 seconds, which accelerates the forecast speed by thousands of times compared to the traditional numerical GOFSs.
@@ 1 Introduction
Ocean forecasting is critically important for many marine activities. At present, the leading GOFSs (e.g. Mercator Ocean Physical SYstem (PSY4) and Real-Time Ocean Forecast System (RTOFS)) use physics-driven models in fluid mechanics and thermodynamics to predict future ocean motion states and phenomena based on current ocean conditions [1]. The GOFSs adopt numerical methods that rely on supercomputers to solve the partial differential equations of the physical models. Due to their desirable performance, they are operationally run in different countries worldwide.
However, numerical forecasting methods are usually computationally expensive and slow. For example, a single forecasting simulation in the numerical GOFSs may take hours on a supercomputer with hundreds of computational nodes [2]. Besides, improving the forecasting accuracy of these methods is exceedingly challenging because they heavily rely on the human cognitive abilities in understanding the physical laws of the ocean environment [3].
@@
    We propose the first data-driven $1/12^{\circ}$ resolution global ocean eddy-resolving forecasting model XiHe, which achieves stronger forecast performance in all testing variables than the compared leading operational numerical GOFSs. Moreover, XiHe can make a forecast in only 0.35 seconds (1000x faster than the numerical GOFSs) averagely with one GPU.
@@ 2.1 Physics-Driven Methods
Until now, numerical ocean forecasting methods have contributed overall the highest prediction accuracy, but they still face two major limitations. Firstly, they are usually computationally expensive and slow due to the high complexity of solving the physical partial differential equations [4]. Secondly, the improvement of the forecasting accuracy is exceedingly challenging due to their limitation of human cognitive abilities in understanding the physical laws of the ocean environment [3].
@@ 3.1 Training Dataset
Considering the impact of the sea surface wind on ocean dynamics, we incorporate the 10m u-component of wind (U10) and 10m v-component of wind (V10) into the training data of XiHe. The wind field data for training is from the Fifth Generation Global Atmospheric Reanalysis data (ERA5) produced by the European Centre for Medium-Range Weather Forecasts (ECMWF) [30]. Their spatial resolution is 0.25° with 137 model levels. In order to standardize the data dimensions, we interpolated the wind data to the same spatial and temporal resolution as the GLORYS12 reanalysis data.
@@ 4.4.2 Global SIE module
${\mathbf{G}^{\prime}}^{l}=\operatorname{Concat}_{\{h\}}\left(\operatorname{Attention}\left(\mathbf{W}_{h}^{Q}\mathbf{G}_{h}^{l},\mathbf{W}_{h}^{K}\mathbf{\tilde{Z}}^{l}_{h},\mathbf{W}_{h}^{V}\mathbf{\tilde{Z}}^{l}_{h}\right)\right),$ (5)
@@ 5 Experimental Evaluation
The forecasting capacity of the XiHe is evaluated based on satellite observations, in situ observations, and the IV-TT Class 4 evaluation framework of the world’s leading operational GOFSs. Note that XiHe forecasts the oceanic variables in a $1/12^{\circ}$ spatial resolution (grid format of 2041$\times$4320), which is the same as the GLORYS12 reanalysis data.
@@ 5.1.1 The IV-TT Class 4 Framework
We evaluated the performances of XiHe under the GODAE OceanView IV-TT Class 4 framework, which is designed to evaluate the world’s leading operational ocean forecasting systems like FOAM, PSY4, GIOPS, BLUElink OceanMAPS, and RTOFS [32]. The IV-TT Class 4 framework focuses on the assessment of the forecast performance and skill, using available ocean observations including in situ drifting buoys and satellite data in near real-time. Within the IV-TT Class 4 framework, each forecasting system interpolates its forecast fields, including ocean temperature profile, ocean salinity profile, SLA, etc., to the observation locations. Then different forecasting systems can be evaluated using the observations and compared with each other. The framework utilizes ocean Temperature and Salinity observations that are sourced from Argo profiles. Additionally, SLA was measured by satellite altimeters [32].
There are five statistical metrics in the IV-TT Class 4 framework, including Bias, root mean square error (RMSE), ACC, and two skill scores constructed using persistence and climatology fields [32]. The RMSE, ACC, and Bias are used in this paper. According to the evaluation of the IV-TT framework, we first interpolate the grid forecast results into the IV-TT observations and then calculate the evaluation metrics [32]. For example, the RMSE and ACC are defined in Eq. (10) and Eq. (11).
@@
${\rm ACC}=\frac{\sum_{i=1}^{N}(\mathbf{F_{i}}-\mathbf{C_{i}})(\mathbf{O_{i}}-\mathbf{C_{i}})}{\sqrt{\sum_{i=1}^{N}(\mathbf{F_{i}}-\mathbf{C_{i}})^{2}}\sqrt{\sum_{i=1}^{N}(\mathbf{O_{i}}-\mathbf{C_{i}})^{2}}},$ (11)
@@
We compare XiHe with the four state-of-the-art global ocean forecasting systems, including PSY4 [17], GIOPS [15], FOAM [16], and BLUElink OceanMAPS (BLK). Details of these forecasting systems are listed in Table I. For RTOFS, considering that there are only four months of forecasting data from Jan. 2019 to Dec. 2020, thus we do not show the experimental results of RTOFS in this paper.
@@ 5.1.2 Overall Evaluation
For all the evaluated variables including ocean temperature, salinity, 15m-depth current, and SLA, XiHe consistently outperforms all the numerical GOFSs. The advantage persists across lead times from 1 day to 10 days. For ocean temperature, compared to FOAM, PSY4, GIOPS, and BLK in the 144-hour (6-day) lead time forecasting, XiHe’s vertical averaged temperature provides an average improvement of 9.94%, 11.10%, 8.94%, and 24.70%, respectively. Meanwhile, XiHe’s vertical averaged salinity provides an average improvement of 5.90%, 9.77%, 11.52%, and 22.98%, respectively.
@@
Compared with the PSY4 system on the 15m-depth zonal and meridional component of ocean currents, XiHe shows a much smaller growing tendency in RMSE. The slopes of the error curves can reflect the growth of forecasting error as the lead time increases. We calculated the approximate slopes of the error curves for PSY4 and XiHe with the lead times from 1 day to 10 days, respectively. For the 15m-depth zonal (meridional) component of ocean currents, the slope of PSY4 is 0.00189 (0.00180) $m/(s\cdot day)$, which is about 3.0 times (4.5 times) as large as 0.00064 (0.00040) $m/(s\cdot day)$ of XiHe. This demonstrates that the forecasting error of XiHe grows much slower than PSY4, and XiHe has the potential for a longer forecast lead time. It is noteworthy that the 10-day RMSE of XiHe is remarkably smaller than the 1-day RMSE of PSY4. Furthermore, to assess the forecasting performance of XiHe for a longer lead time, we train the models for the lead time of 20 days, 30 days, and 60 days. As shown in Fig 3, the RMSE values of the 15m-depth zonal component of ocean currents (uo) for the three lead times are 0.2038, 0.2093, and 0.2165 respectively, which is even smaller than the PSY4’s 10-day RMSE of 0.2214. The RMSE values of the 15m-depth meridional component of ocean currents (vo) for the three lead times are 0.1955, 0.2008, and 0.2062 respectively, which is smaller than the 10-day RMSE of 0.2149 for PSY4. This indicates that the accuracy of ocean current forecasting of XiHe out to 60 days is even better than that of PSY4 in 10 days.
@@
From the perspective of both the oceanic variables and the lead times, the overall evaluation results firmly demonstrate that the data-driven XiHe model has achieved stronger forecast performance than the leading operational numerical GOFSs including PSY4, GIOPS, BLUElink OceanMAPS, and FOAM.
@@
Figure 4: The performance comparison of our data-driven model XiHe with the four leading operational numerical GOFSs based on the IV-TT framework. The x-axis represents the forecast time. The y-axis represents the forecasting ACC (higher is better) of the global ocean forecasting systems. In addition, for each lead time, the ACC value is averaged from Jan. 1st, 2019 to Dec. 31st, 2020.
@@
As shown in Fig. 4 of the ACC analysis, XiHe outperforms the four leading operational numerical GOFSs in terms of temperature and salinity. The ACC values of temperature and salinity profiles of XiHe’s 20-day forecasting are 0.8743 and 0.6420, which are even better than the 10-day forecasts of other GOFSs. This highlights XiHe’s robust forecasting capabilities.
@@
To be more detailed, we present the daily RMSE of temperature, salinity, uo, vo, and SLA from Jan. 1st, 2019 to Dec. 31st, 2020 in Fig. 6 with the lead times of 1-day, 5-day, and 10-day. For the five variables, XiHe exceeds the numerical forecasting systems by an immense margin during the evaluated period. Especially for the uo and vo variables, XiHe achieves an average improvement of 10.16% and 11.21% over PSY4 in the 10-day lead time forecasting, respectively. It can be observed that the forecast errors for different variables exhibit certain seasonal variation characteristics. For example, with regard to the XiHe and PSY4 models, the forecast errors for ocean currents (uo, vo) are smaller in the boreal summer and larger in the boreal winter. As the forecast lead time increases, the seasonal variation in forecast errors becomes more evident. In addition, it is worth mentioning that the forecast results provided by XiHe are more stable than PSY4, exhibiting fewer excessive peak values that would indicate dramatic errors.
@@ 5.1.3 Temperature/Salinity Profiles Evaluation
In order to gain more insights into the forecast errors at different depths, Fig. 7 and Fig. 8 show the profiles of the temperature and the salinity RMSE for different GOFSs in depth above 650m.
As to the depth-wise evaluation results towards the salinity variable in Fig. 7, XiHe outperforms all the compared numerical operational GOFSs almost in the whole depths. And only when the lead time is 1 day and 5 days, for the salinity profiles just from 95m to 137.5m, can FOAM, GIOPS and PSY4 achieve comparable or slightly better performances than XiHe. This denotes that XiHe performs overall better than the leading GOFSs in accurately forecasting the salinity variable across the entire profile.
The experimental results in Fig. 8 show the depth-wise evaluation of temperature. XiHe consistently outperforms the four operational numerical GOFSs in terms of temperature variables almost in the whole depths, which is particularly excellent below the depth of 30m. And only when the lead time is 1 day and 5 days, for the temperature profiles just above 30m, can FOAM and GIOPS and achieve slightly better performances than XiHe. This denotes that XiHe performs overall better than the leading GOFSs in accurately forecasting the temperature variable across the entire profile.
Figure 7: The profiles of salinity RMSE above 650m for different GOFSs. The lead times are 1 day, 5 days and 10 days. For each lead time, the reported RMSE is averaged from Jan. 1st, 2019 to Dec. 31st, 2020.
Figure 8: The profiles of temperature RMSE above 650m for different GOFSs. The lead times are 1 day, 5 days and 10 days. For each lead time, the reported RMSE is averaged from Jan. 1st, 2019 to Dec. 31st, 2020.